OpenStudy (anonymous):

how do you find a particle's minimum speed when given a distance formula s(t)

OpenStudy (anonymous):

first you have to find the formula for speed which is distance over time. Then you can set time to 0 which will give you your minimum speed

OpenStudy (anonymous):

the formla is s(t)=t^3-12t^2+45t-17

OpenStudy (anonymous):

i'm not given distances, or a time interval

OpenStudy (anonymous):

based on the equation , the question is , does hte particle have a minimum speed on the interval t\[\ge\]0, and if so what is it and at what time is it attained

OpenStudy (anonymous):

velocity or speed is the derivative of the position function so s'(t)=3t^2-24t+45 now just set t=0. So, s'(0)=3(0)^2-24(0)+45=45